# Interpretable Embeddings of Speech Enhance and Explain Brain Encoding Performance of Audio Models

### v1

@@ Abstract
Self-supervised speech models (SSMs) are increasingly hailed as more powerful computational models of human speech perception than models based on traditional hand-crafted features. However, since their representations are inherently black-box, it remains unclear what drives their alignment with brain responses. To remedy this, we built linear encoding models from six interpretable feature families: mel-spectrogram, Gabor filter bank features, speech presence, phonetic, syntactic, and semantic Question-Answering features, and contextualized embeddings from three state-of-the-art SSMs (Whisper, HuBERT, WavLM), quantifying the shared and unique neural variance captured by each feature class. Contrary to prevailing assumptions, our interpretable model predicted electrocorticography (ECoG) responses to speech more accurately than any SSM. Moreover, augmenting SSM representations with interpretable features yielded the best overall neural predictions, significantly outperforming either class alone.
Further variance-partitioning analyses revealed previously unresolved components of SSM representations that contribute to their neural alignment: 1. Despite the common assumption that later layers of SSMs discard low-level acoustic information, these models compress and preferentially retain frequency bands critical for neural encoding of speech (100–1000 Hz). 2. Contrary to previous claims, SSMs encode brain-relevant semantic information that cannot be reduced to lower-level features, improving with context length and model size. These results highlight the importance of using refined, interpretable features in understanding speech perception.
@@ 1 Introduction
Self-supervised models have emerged as increasingly popular tools for studying how the brain processes language and speech. Numerous studies have employed large language models (LLMs) and demonstrated their effectiveness in predicting neural responses to narrative stimuli by linearly probing their internal representations (Toneva and Wehbe 2019; Goldstein et al. 2022; Caucheteux and King 2022; Antonello, Vaidya, and Huth 2023; Hosseini et al. 2024; Mischler et al. 2024). Similarly, self-supervised speech models (SSMs) have been used to investigate neural representations underlying speech perception. These studies commonly conclude that, like LLMs, SSMs are highly effective in predicting brain responses to speech, often outperforming their interpretable baselines (Vaidya, Jain, and Huth 2022; Anderson, Davis, and Lalor 2024; Goldstein et al. 2025; Tuckute et al. 2023). Consequently, these findings are interpreted as evidence that SSMs are more powerful and brain-like models of speech perception, with their superiority often attributed to their ability to integrate contextual information over time (Anderson, Davis, and Lalor 2024; Goldstein et al. 2025).
However, despite the apparent success of SSMs in predicting brain responses, several critical challenges remain in drawing meaningful neuroscientific conclusions from these results. First, the goal of cognitive neuroscience is not to build computational models with the best predictive power, but to understand the mechanisms underlying such neural computation. While there has been growing effort to investigate what these models encode, much of the existing work has focused on characterizing their internal representations (Pasad, Chou, and Livescu 2022; Pasad, Shi, and Livescu 2023; Pasad et al. 2024; Ashihara et al. 2023; Shen et al. 2023; Choi and Yeo 2022; Choi et al. 2024), an important but distinct goal from identifying which aspects of their representations are relevant for neural processing. Techniques such as Canonical Correlation Analysis (CCA), Representational Similarity Analysis (RSA), and linear probing with interpretable embeddings are commonly used to probe the internal structure of SSMs. However, these methods typically assess aggregate similarity across representational spaces and do not isolate specific components that are mechanistically meaningful for brain function. As a result, even when combined with strong neural prediction performance, such findings offer limited guidance for understanding how model representations map onto neural processes and may be prone to misinterpretation. Second, the predictive performance of SSMs may potentially be explained by missed, yet simple confounds. In the context of LLMs, Hadidi et al. showed that most of the neural variance explained by GPT2-XL can be explained by simple confounds such as word-rate and positional information (Hadidi et al. 2025). This raises the concern that SSMs, too, might be over-credited with encoding unique brain-relevant contextual information when, in fact, their performance is driven by simpler, interpretable confounds. Collectively, these challenges underscore the importance of disentangling what properties drive the predictive performance of SSMs and assessing the extent to which their contextual representations capture unique neural variance beyond a broad set of interpretable features.
A detailed examination of the existing empirical evidence further highlights these concerns. While there is robust evidence that SSMs yield better alignment in higher-order brain regions than simple acoustic baselines (Vaidya, Jain, and Huth 2022; Li et al. 2023), the nature of the representations driving this alignment remains unclear, particularly with regard to whether they encode brain-relevant semantic information. For example, Anderson et al. (Anderson, Davis, and Lalor 2024) found that combining representations from Whisper’s later layers with those of GPT-2 did not improve alignment with EEG data compared to Whisper alone, suggesting that Whisper already encodes lexical content relevant for neural processing. In contrast, Oota et al. (Oota et al. 2024) found that after regressing out low-level acoustic and speech features, the representations of the SSMs showed no significant alignment with high-level language regions of the brain, indicating that these models may lack higher-order semantic information. We argue that these contrasting findings may stem from the indirect nature of current methodologies, which often involve inferring properties of one black box (SSMs) using another (the brain) without the use of interpretable controls, making interpretation inherently challenging.
In contrast to previous work, this study directly addresses these open questions using a variance partitioning approach applied to hand-crafted, interpretable features spanning low-level acoustic to high-level lexical domains. Specifically, we show that: 1. Handcrafted interpretable features can outperform SSM representations in predicting brain responses. Combining interpretable features and SSM representations leads to the best performance. 2. SSM representations retain brain-relevant low-level acoustic information across layers, with acoustic features becoming increasingly compact to accommodate higher-level representations. 3. SSMs encode semantic information that is relevant to neural processing, which improves with context length and model size.
Figure 1: Summary of the methodologies. (a) This study constructs six sets of interpretable features—mel-spectrograms, Gabor filter bank features, speech presence, phonetic, syntactic, and semantic features—and compares them to representations from three state-of-the-art self-supervised speech models (SSMs): Whisper, HuBERT, and WavLM. The comparison is based on each representation’s ability to predict neural responses, measured through Pearson’s $r$ and variance explained $R^{2}$. (b) An example of each linguistic feature set. The word is represented as phonetic (ARPAbet), syntactic (Part-of-Speech Tag, Dependency Parsing, stop word or not), and semantic (Question Answering) features. (c) Diagram of the variance partitioning approach.
@@ 2 Methods
Fig. 1 presents a summary of the methodologies employed in this research. All results can be reproduced using the code available at: https://github.com/rikishimizu/interpretable˙speech
@@ 2.1 Datasets
All results were derived using open-source data contained in the Podcast ECoG Dataset (Zada et al. 2025). It contains electrocorticography (ECoG) recordings from nine patients with epilepsy who listened to a 30-minute excerpt of the podcast This American Life (“So a Monkey and a Horse Walk Into a Bar: Act One—Monkey in the Middle”). Of the 1,330 intracranial electrodes originally implanted, 1,268 were retained after excluding contacts with imprecise anatomical localization or excessive noise in their power-spectral density. Electrode positions were projected onto a template cortical surface for visualization. Full pre-processing details are provided in (Zada et al. 2025).
@@
We focused on broadband high-gamma activity (70–200 Hz), extracted with a fourth-order zero-phase Butterworth band-pass IIR filter and converted to analytic amplitude. The resulting envelopes were down-sampled to 4 Hz and z-scored.
2.2 Interpretable auditory features
@@
We derived six families of interpretable predictors: (i) mel-spectrograms, (ii) Gabor filter bank (GBFB) features, (iii) speech presence, (iv) phonetic features, (v) semantic features, and (vi) syntactic features (Fig. 1a). These features reflect the language processing hierarchy in speech comprehension (Heer et al. 2017; Keshishian et al. 2023). For phonetic, semantic, and syntactic features, the feature vector at time $t$ s was computed by summing up the embeddings for all words or phonemes whose offset time lies in [$t-0.5$, $t$]s. Full details of each feature processing are available in Section S2.
Mel-spectrogram
@@
We computed mel-spectrograms from 0–8 kHz with 32 mel bins and further converted them to decibels using librosa (McFee et al. 2015).
Gabor bank filter features (GBFB)
@@
Spectro-temporal modulation features were extracted with the Gabor filter-bank implementation of Schädler et al (Schädler, Meyer, and Kollmeier 2012).
Speech presence
@@
Speech probability was estimated with EfficientAT, a CNN-based audio tagger (Schmid, Koutini, and Widmer 2023).
Phonetic features
@@
Each word in the transcript was converted to its ARPAbet phoneme representations using nltk.corpus.cmudict (Bird and Loper 2004). Every phoneme was one-hot encoded over the 39 ARPAbet symbols.
Semantic features
Following Benara et al. 2024, we obtained 168-dimensional question-answering (QA) semantic embeddings by prompting GPT-4 (gpt-4-0125-preview; OpenAI 2023) with 56 yes/no questions about each word’s meaning (e.g., “Does the word/sentence involve an expression of personal values or beliefs?”) with 3 different context lengths (word-level, 1.5 s, and 3 s). For the 1.5/3-second context length, the model was additionally provided with the words spoken in the 1.5/3 seconds preceding each target word when answering the questions. Because of how this feature is constructed, its effective context window can extend up to approximately 4.75 seconds—accounting for the 3-second input context, the additional 0.5-second summing window, and the maximum word duration (about 1.25 seconds for words like “approximately”). A full list of questions is provided in Table S2.
Syntactic features
@@
Following Zada et al. 2025, each word was first converted into 96-dimensional embeddings, comprising 50 part-of-speech tags, 45 dependency labels, and a stop/non-stop label using the spaCy (Honnibal et al. 2020).
2.3 Speech Audio Models
We used three speech models: Whisper (Radford et al. 2022), HuBERT (Hsu et al. 2021), and WavLM (Chen et al. 2022). (Fig. 1a)
@@ 2.3 Speech Audio Models
For this work, unless otherwise specified, we used the largest variants of these models: Whisper Large V1 (638M Parameter, 32 layers), HuBERT X-large (964M Parameter, 48 layers), and WavLM X-large (317M Parameter, 24 layers). Table S1 shows the summary of the architecture and training objectives for each model. To generate the embeddings with a context length of $Ctx$ seconds, we applied a sliding window of $Ctx$ seconds and a stride of 250 ms to the input audio. For each window, the representation at a given layer was defined as the hidden state of the final token. For Whisper, we only used hidden states from the encoder. In this study, $Ctx$ was varied across the following values: 0.5, 1, 2.5, 5, 10, 20, and 30 seconds. PCA with $95\%$ variance was applied to each model’s embeddings to reduce dimensionality.
@@ 2.4 Encoding Model Construction
After processing, the high gamma signals we predict have shape $[T,N_{ch}]$ where $T$ is the number of time points (i.e., 7200) and $N_{ch}$ is the number of channels (i.e., 1268). To estimate the model performance, we used a 5-fold cross-validation scheme, splitting the signals into 5 temporally contiguous chunks. For each fold, the model was trained on $80\%$ of the data and tested on the remaining $20\%$. Within the training folds, we employed a 4-fold nested cross-validation scheme to optimize the model parameters. Unless otherwise noted—specifically in analyses examining the impact of context size on encoding performance—the optimal context size was selected and used for each speech audio model.
Banded ridge regression
@@
To create a joint encoding model with features from different feature spaces and dimensionalities, we employed banded ridge regression, which has been shown to improve neural encoding performance when using diverse feature spaces (Nunez-Elizalde, Huth, and Gallant 2019). Unlike ridge regression with spherical Gaussian priors, banded ridge regression optimizes the regularization parameter alpha separately for each feature space. We used the himalaya library by Dupre La Tour et al. (Dupré la Tour et al. 2022). Each feature was z-scored before applying the regression model. Unless otherwise specified, we varied alpha over 20 logarithmically spaced points between $10^{0}$ and $10^{3}$.
Performance Metrics

### v2

# Interpretable Embeddings of Speech Explain and Enhance the Brain Encoding Performance of Audio Models

@@ Abstract
Speech foundation models (SFMs) are increasingly hailed as powerful computational models of human speech perception. However, since their representations are inherently black-box, it remains unclear what drives their alignment with brain responses. To remedy this, we built linear encoding models from six interpretable feature families: mel-spectrogram, Gabor filter bank features, speech presence, phonetic, syntactic, and semantic features, and contextualized embeddings from three state-of-the-art SFMs (Whisper, HuBERT, WavLM), quantifying electrocorticography (ECoG) response variance shared between feature classes. Variance-partitioning analyses revealed several key insights: First, the SFMs’ alignment with the brain can be mostly explained by their ability to learn and encode simple interpretable speech features. Second, SFMs exhibit a systematic trade-off between encoding of brain-relevant low-level and high-level features across layers. Finally, our results show that SFMs learn brain-relevant semantics which cannot be explained by lower-level speech features, with this capacity increasing with model size and context length. Together, our findings suggest a principled approach to build more interpretable, accurate, and efficient encoding models of the brain by augmenting SFM embeddings with interpretable features.
Disclosure: Portions of this manuscript were assisted by OpenAI ChatGPT for writing. The authors reviewed, edited, and take full responsibility for the content.
@@ 1 Introduction
Large-scale foundation models have become increasingly popular tools for studying how the brain processes language and speech. Trained on vast amounts of linguistic and acoustic data, these models appear to extract rich features relevant for predicting neural activity. Numerous studies have employed large language models (LLMs) and demonstrated their effectiveness in predicting neural responses to narrative stimuli by constructing linear models based on their internal representations (Toneva & Wehbe, 2019; Goldstein et al., 2022; Caucheteux & King, 2022; Antonello et al., 2023; Hosseini et al., 2024; Mischler et al., 2024). Similarly, speech foundation models (SFMs) have been used to investigate neural representations underlying speech perception. These studies commonly conclude that, like LLMs, SFMs are highly effective in predicting brain responses to speech due to their ability to integrate contextual information over time (Vaidya et al., 2022; Millet et al., 2022; Anderson et al., 2024; Goldstein et al., 2025; Tuckute et al., 2023; Antonello et al., 2023).
Despite these promising results, the neuroscientific basis of SFM’s predictive performance remains to be thoroughly investigated. The existing work on SFM-brain alignment has largely focused on SFM’s predictive success, while offering limited insight into which specific representational components drive their alignment with the brain. In parallel, complementary work in machine learning has examined SFM’s internal representations using probing and representational similarity analysis (Pasad et al., 2022; 2023; 2024; Ashihara et al., 2023; Shen et al., 2023; Choi & Yeo, 2022; Choi et al., 2024). Results from this line of work indicate that SFM representations encode a wide range of features: early layers often reflect acoustic and phonetic properties, whereas later layers increasingly gain high-level phonological, morphological, syntactic, and semantic information. While these studies have revealed important aspects of SFM representations, they do not test which specific representational components actually drive SFMs’ brain encoding performance, and thus leave a critical gap in our mechanistic understanding of SFM-brain alignment.
Due to this gap, several key questions regarding the nature of SFM-brain alignment remain unanswered. For example, several studies report an almost monotonic increase in brain encoding performance across SFM layers, with the latest layers in the network substantially outperforming earlier layers (Vaidya et al., 2022; Antonello et al., 2023; Anderson et al., 2024). A key open question is why this trend occurs. Does it reflect a genuine accumulation of increasingly high-level, brain-relevant information, with low-level details retained to support neural encoding? Or does it mask a trade-off, where gains in high-level representations come at the expense of lower-level features? Even more fundamentally, although the main representational advantage of SFMs is that they are not constrained to an interpretable feature space, it remains unclear how much benefit this provides beyond simpler interpretable features. The central question, then, is whether SFMs’ brain encoding performance primarily reflects their capacity to capture speech information in highly non-linear and uninterpretable ways, or instead derives largely from their ability to learn and encode simpler, interpretable features.
These unresolved questions highlight the need to identify which specific features in SFM representations underlie their capacity to predict brain activity at different layers, and to determine whether their contextual representations explain unique neural variance beyond that which can be explained by interpretable features alone. This study directly addresses these open questions using intracranial electrocorticography (ECoG), which has high temporal and spatial resolution that makes it suitable for studying both low-level and high-level feature encoding. We analyze this data with a variance partitioning approach applied to hand-crafted, interpretable features spanning acoustic to lexical domains. Our analysis reveals: First, the SFMs’ alignment with the brain can be mostly explained by their ability to learn and encode simple, interpretable speech features. Second, SFMs exhibit a systematic trade-off between encoding of brain-relevant low-level and high-level features across layers. Finally, our results show compellingly that SFMs learn brain-relevant semantics that cannot be explained by lower-level speech features, with this capacity increasing with model size and context length. Together, these findings suggest a principled approach to building more interpretable, accurate, and efficient encoding models of the brain by augmenting SFM embeddings with interpretable features.
@@ 2 Methods
Fig.˜1 presents a summary of the methodology employed in this research.
@@ 2.1 Datasets
All results were derived using open-source data contained in the Podcast ECoG Dataset (Zada et al., 2025). It contains electrocorticography (ECoG) recordings from nine patients with epilepsy who listened to a 30-minute excerpt of the podcast This American Life (“So a Monkey and a Horse Walk Into a Bar: Act One—Monkey in the Middle”). Of the 1,330 intracranial electrodes originally implanted, 1,268 were retained after excluding contacts with imprecise anatomical localization or excessive noise in their power-spectral density. Electrode positions were projected onto a template cortical surface for visualization. Full pre-processing details are provided in Zada et al. (2025).
@@
We focused on broadband high-gamma activity (70–200 Hz), extracted with a fourth-order zero-phase Butterworth band-pass Infinite Impulse Response (IIR) filter and converted to analytic amplitude. The resulting envelopes were down-sampled to 4 Hz and z-scored.
Figure 1: Summary of the methodology. (a) We constructed six sets of interpretable features—mel-spectrograms, Gabor filter bank features, speech presence, phonetic, syntactic, and semantic features—and compared them to representations from three state-of-the-art speech foundation models (SFMs): Whisper, HuBERT, and WavLM. The comparison is based on each representation’s ability to predict neural responses, measured through Pearson’s $r$ and the variance explained $R^{2}$. (b) An example of each linguistic feature set. The word is represented as phonetic (ARPAbet), syntactic (Part-of-Speech Tag, Dependency Parsing, stop word or not), and semantic (Question Answering) features. (c) Diagram of the variance partitioning approach.
2.2 Interpretable features
@@
We derived six families of interpretable predictors: (i) mel-spectrograms, (ii) Gabor filter bank (GBFB) features, (iii) speech presence, (iv) phonetic features, (v) syntactic features, and (vi) semantic features (Fig.˜1a). These features reflect the language processing hierarchy in speech comprehension (Heer et al., 2017; Keshishian et al., 2023). As shown in Fig.˜S1, each feature explains unique neural variance $R^{2}$ relative to the others, demonstrating that they capture complementary aspects of brain responses to speech.
@@
(i) Mel-spectrograms. We computed mel-spectrograms from 0–8 kHz with 32 mel bins and further converted them to decibels using librosa (McFee et al., 2015). To create a feature vector for time $t$ s, we concatenated the frames spanning [t0.5t-0.5, $t$]s, yielding a 672-dimensional vector. Dimensionality was reduced by principal-component analysis (PCA), retaining 95% of the variance.
@@
(ii) Gabor filter bank (GBFB) features. Spectro-temporal modulation features were extracted with the Gabor filter-bank implementation of Schädler et al. (2012). The 455-dimensional features, computed at 100 Hz, were down-sampled to 4 Hz to align with the neural data; PCA (95% variance) was applied prior to modeling.
@@
(iii) Speech presence. Speech probability was estimated with EfficientAT, a CNN-based audio tagger (Schmid et al., 2023). To produce a feature scalar for time $t$ s, we fed the model the audio spanning [t0.5t-0.5, $t$]s.
@@
(iv) Phonetic features. Each word in the transcript was converted to its ARPAbet phoneme representations using nltk.corpus.cmudict (Bird & Loper, 2004). Word intervals were evenly subdivided among their $N$ phones, assigning on- and offset times to every phoneme. Every phoneme was one-hot encoded over the 39 ARPAbet symbols. To create a feature vector for time $t$ s, the embeddings for every phoneme whose offset time lies in [t0.5t-0.5, $t$]s were summed, producing a $T\times 39$ matrix.
@@
(v) Syntactic features. Following Zada et al. (2025), each word was first converted into 96-dimensional embeddings, comprising 50 part-of-speech tags, 45 dependency labels, and a stop/non-stop label using the spaCy (Honnibal et al., 2020). To create a feature vector for time $t$ s, the embeddings for every word whose offset time lies in [t0.5t-0.5, $t$]s were summed, producing a $T\times 96$ matrix.
(vi) Semantic features. Following Benara et al. (2024), we obtained 168-dimensional question-answering (QA) semantic embeddings by prompting GPT-4 (gpt-4-0125-preview; OpenAI (2023)) with 56 yes/no questions about each word’s meaning (e.g., “Does the word/sentence involve an expression of personal values or beliefs?”) with 3 different context lengths (word-level, 1.5 s, and 3 s). For the 1.5/3-second context length, the model was additionally provided with the words spoken in the 1.5/3 seconds preceding each target word when answering the questions. To create a feature vector for time $t$ s, the embeddings for every word whose offset time lies in [t0.5t-0.5, $t$]s were summed, producing a $T\times 168$ matrix. Because of how this feature is constructed, its effective context window can extend up to approximately 4.75 seconds—accounting for the 3-second input context, the additional 0.5-second summing window, and the maximum word duration (about 1.25 seconds for words like “approximately”). A full list of questions is provided in Table˜S2.
2.3 Speech Models
@@
For this work, unless otherwise specified, we used the largest variants of three speech models: Whisper Large V1 (638M Parameters, 32 layers) (Radford et al., 2022), HuBERT X-large (964M Parameters, 48 layers) (Hsu et al., 2021), and WavLM Large (317M Parameters, 24 layers) (Chen et al., 2022). This choice was motivated by prior findings from Antonello et al. (2023), which showed that the brain encoding performance of SFM models increased logarithmically with the number of model parameters. Table˜S1 shows the summary of the architecture and training objectives for each model. To generate the embeddings with a context length of $Ctx$ seconds, we applied a sliding window of $Ctx$ seconds and a stride of 250 ms to the input audio. For each window, the representation at a given layer was defined as the hidden state of the final token. We made this choice to remain consistent with prior work (Vaidya et al., 2022; Antonello et al., 2023; Oota et al., 2024) and to maximize interpretability, since the encoding results depend on a single causal token from each layer, making them easier to compare and interpret. For Whisper, we only used hidden states from the encoder. In this study, $Ctx$ was varied across the following values: 0.5, 1, 2.5, 5, 10, 20, and 30 seconds. PCA with $95\%$ variance was applied to each model’s embeddings to reduce dimensionality.
@@ 2.4 Encoding Model Construction
After processing, the high gamma signals we predict have shape $T,N_{ch}]$ where $T$ is the number of time points (i.e., 7200) and $N_{ch}$ is the number of channels (i.e., 1268). To estimate the model performance, we used a 5-fold cross-validation scheme, splitting the signals into 5 temporally contiguous chunks. For each fold, the model was trained on $80\%$ of the data and tested on the remaining $20\%$. Within the training folds, we employed a 4-fold nested cross-validation scheme to optimize the model parameters.
Prior to conducting variance partitioning, we computed encoding performance values from embeddings with different context lengths for each model (Fig.˜S2). The context lengths yielding the highest test correlations across layers were 30 seconds for HuBERT, 5 seconds for WavLM, and 30 seconds for Whisper. Unless otherwise noted, all subsequent analyses are based on embeddings with these optimal context lengths.
2.4.1 Banded ridge regression
@@
To create a joint encoding model with features from different feature spaces and dimensionalities, we employed banded ridge regression, which has been shown to improve neural encoding performance when using diverse feature spaces (Nunez-Elizalde et al., 2019). Unlike ridge regression with spherical Gaussian priors, banded ridge regression optimizes the regularization parameter separately for each feature space. We used the himalaya library by Dupré la Tour et al. (2022). The regression formulation becomes:
$\bm{\beta}^{\star}=\arg\min_{\bm{\beta}}\left\|\sum_{i=1}^{m}\mathbf{X}_{i}\bm{\beta}_{i}-\mathbf{y}\right\|_{2}^{2}\;+\;\sum_{i=1}^{m}\alpha_{i}\left\|\bm{\beta}_{i}\right\|_{2}^{2}.$
Here, $\mathbf{X}_{i}\in\mathbb{R}^{T\times p_{i}}$ is the feature matrix for the $i$-th feature space, $\bm{\beta}_{i}\in\mathbb{R}^{p_{i}}$ is the corresponding vector of regression weights, $\mathbf{y}\in\mathbb{R}^{T}$ is the observed response vector (e.g., neural responses), $\alpha_{i}\in\mathbb{R}_{\geq 0}$ is the scalar regularization parameter for the $i$-th feature space, and $m$ is the (scalar) number of feature spaces. The optimal coefficient vector concatenated across spaces is $\bm{\beta}^{\star}=\big[\bm{\beta}_{1}^{\top},\dots,\bm{\beta}_{m}^{\top}\big]^{\top}\in\mathbb{R}^{p}$ with $p=\sum_{i=1}^{m}p_{i}$.
Each feature was z-scored before applying the regression model. Unless otherwise specified, we varied alpha over 20 logarithmically spaced points between $10^{0}$ and $10^{3}$.
2.4.2 Performance Metrics